1). Given a marginal propensity to consume = .8

(Ceteris paribus) and the government increases the level of transfer payments by $100,

we should expect that the GDP will increase by

a. $500.

b. $80.

c. $100.

d. $400.

2). Suppose that actual GDP (Y) for France is 100 euros and:

Consumption = 10 + .5Y

Investment Spending = 5

Government Expenditure = 20

The Marginal Tax Rate (t) = .20

At these current levels, the size of the budget deficit is

a. 10.

b. 35.

c. 20.

d. 0.

3). The most appropriate fiscal policy for stimulating economic growth and thereby reducing cyclical unemployment would be

a. to raise the natural rate of unemployment.

b. to increase the value of the domestic currency in foreign exchange.

c. increasing the money supply.

d. increasing transfer payments to the population.

4). GDP (Y) = C + I + G

Consumption = 10 + .75Y

Investment = 20 + .15Y

Government = 35

The size of the expenditure multiplier for this hypothetical economy is

a. 4.

b. 10.

c. none of the other answers are correct.

d. 2.
Answer #1

1. d. $400.
(According to the multiplier formula,
Change in GDP/Change in transfer payments = MPC/(1-MPC) = .8/(1-.8) = .8/.2 = 4
So, change in GDP = Change in transfer payments*4 = 100*4 = $400)

2. d. 0.
(Budget deficit = Government expenditure - Tax = 20 - .20*(100) = 20 - 20 = 0)

3. d. increasing transfer payments to the population.
(Fiscal policy includes increasing the transfer payments which will stimulate economic growth.)

4. b. 10.
(MPC = .75; MPI = .15
Expenditure multiplier = 1/(1-MPC-MPI) = 1/(1-.75-.15) = 1/.1 = 10)